# Gyro-based Neural Single Image Deblurring

### v2

@@ Abstract
In this paper, we present GyroDeblurNet, a novel single image deblurring method that utilizes a gyro sensor to effectively resolve the ill-posedness of image deblurring. The gyro sensor provides valuable information about camera motion during exposure time that can significantly improve deblurring quality. However, effectively exploiting real-world gyro data is challenging due to significant errors from various sources including sensor noise, the disparity between the positions of a camera module and a gyro sensor, the absence of translational motion information, and moving objects whose motions cannot be captured by a gyro sensor. To handle gyro error, GyroDeblurNet is equipped with two novel neural network blocks: a gyro refinement block and a gyro deblurring block. The gyro refinement block refines the error-ridden gyro data using the blur information from the input image. On the other hand, the gyro deblurring block removes blur from the input image using the refined gyro data and further compensates for gyro error by leveraging the blur information from the input image. For training a neural network with erroneous gyro data, we propose a training strategy based on the curriculum learning. We also introduce a novel gyro data embedding scheme to represent real-world intricate camera shakes. Finally, we present a synthetic dataset and a real dataset for the training and evaluation of gyro-based single image deblurring. Our experiments demonstrate that our approach achieves state-of-the-art deblurring quality by effectively utilizing erroneous gyro data.
@@ 1 Introduction
Blur caused by camera shakes severely degrades image quality as well as the performance of various computer vision tasks such as classification, object detection, and semantic segmentation. To overcome the image degradation caused by blur, single image deblurring has been extensively studied for decades [6, 12, 52, 32, 48, 17, 14]. Recently, various DNN-based approaches have been proposed and have shown significant performance improvements over classical approaches [31, 53, 44, 23, 24, 51, 5, 4]. Nonetheless, they still fail on images with large blur due to the severe ill-posedness of deblurring, i.e., multiple different blur trajectories could lead to similar-looking blurry images.
@@
To address the ill-posedness, several attempts have been made to exploit gyro sensors with which most smartphones are nowadays equipped. As gyro data are correlated with motion of the camera mounted on a smartphone, classical approaches based on blur models use gyro data to guide blur kernel estimation [17, 40, 41, 11, 14, 29]. However, their performances are limited by their restrictive blur models that cannot effectively handle noise, nonlinear camera response functions, saturated pixels, moving objects, and so on. With the advance of deep learning, DNN-based approaches have also been proposed to exploit gyro sensors [30, 25, 16]. Thanks to the remarkable capability of DNNs and gyro sensors, they show better performance compared to classical gyro-based approaches and recent DNN-based approaches without gyro sensors.
@@
However, recent DNN-based approaches that utilize gyro data are still limited in handling real-world blurred images. These approaches assume that the gyro data can accurately represent the blur in the blurred image, but this assumption does not hold in practice. First, gyro data from mobile devices such as smartphones usually have a significant amount of noise caused by various sources [19, 10, 17]. Second, gyro sensors measure the angular velocity around their center. Thus, to exploit gyro sensors, the recent gyro-based methods assume that the gyro sensor and the camera share the same center, and the camera is shaken only by rotational motions. However, in practice, the positions of the camera and the gyro sensor are different, and the camera shakes are caused by more complex motions.Third, real-world blurred images may have moving objects with different blur trajectories that cannot be captured by gyro sensors. All these factors make the motion information encoded in the gyro data inconsistent with the blur in the blurred image, and make it difficult to exploit the gyro data in real-world image deblurring. In this paper, we refer to such inconsistency between the gyro data and blur as gyro error.
@@
In this paper, we propose GyroDeblurNet, a novel gyro-based single image deblurring approach that can produce high-quality deblurring results in the presence of gyro error. To address the gyro error, GyroDeblurNet adopts a carefully designed network architecture that contains a novel gyro refinement block and a novel gyro deblurring block. The gyro refinement block refines gyro data with error using the blur information in the input image. On the other hand, the gyro deblurring block removes blur in the input image using the refined gyro data. To further compensate for gyro error, the gyro deblurring block also exploits the blur information in the input image as well as the input gyro data. Despite the proposed blocks, training a neural network with erroneous gyro data is difficult as the network could be trained to simply ignore the input gyro data. To resolve this issue, we also propose a curriculum-learning-based training strategy [2].
@@
To handle complex real-world camera shakes, we also present the camera motion field, a novel gyro data embedding scheme that can represent complex camera motions. Real-world blurry images often exhibit complex blur trajectories due to random camera shakes, which often occur when shooting a still shot with a long exposure time. Nonetheless, recent gyro-based approaches [30, 25, 16] assume smooth camera motions and represent camera motions using one or two motion vector per pixel or a couple of homographies for the entire image, which leads to low-quality deblurring results. To resolve this issue, our camera motion field is designed to represent complex camera motions using multiple vectors per pixel while maintaining a comparable memory footprint exploiting the spatially smooth nature of camera shakes.
@@
For the training and evaluation of gyro-based deblurring, previous methods [30, 25, 16] rely on synthetic datasets. However, these datasets do not accurately reflect the noise and blur in real-world blurred images as shown by Rim et al. [35]. Additionally, their gyro data do not reflect real-world camera shakes that occur when capturing still-shot images either. Gyro data of previous methods were generated by random sampling [54] or obtained from the Visual-Inertial dataset [38]. However, the Visual-Inertial dataset is originally developed for visual odometry and SLAM, so its camera motions are collected from constantly moving cameras, which differ from camera shakes of still-shot images.
@@
For training and evaluation of gyro-based deblurring for real-world blurred images, we propose two datasets: GyroBlur-Synth and GyroBlur-Real. GyroBlur-Synth is a synthetic dataset for both training and evaluation. The dataset consists of sharp and blurred image pairs with their corresponding gyro data. The gyro data are acquired from a smartphone device to reflect the real-world camera shakes. The blurred images are synthetically generated using the acquired gyro data and the blur synthesis process of RSBlur [34] for realistic blur. GyroBlur-Real is a real-world blur dataset for qualitative evaluation with no ground-truth sharp images. GyroBlur-Real provides 100 real-world blurred images paired with gyro data collected from a smartphone.
@@
  • We propose a novel gyro-based single image deblurring approach, GyroDeblurNet. Extensive quantitative and qualitative evaluations show that GyroDeblurNet outperforms existing single image deblurring methods.
  • To handle gyro error, we introduce a carefully designed network equipped with a novel gyro refinement block and a novel gyro deblurring block. We also develop a novel gyro data embedding scheme to represent complex camera shakes and propose a curriculum learning-based training scheme.
@@ 2 Related Work
To resolve the ill-posedness of deblurring, there have been attempts to leverage inertial measurement sensors such as the gyro sensor and accelerometer as they provide valuable information on the rotational and translational motion of a camera, respectively. Joshi et al. [17] and Hu et al. [14] exploit the gyro sensor and accelerometer to estimate spatially-varying blur kernels. On the other hand, Park and Levoy [11] and Mustaniemi et al. [29] use only the gyro sensor to estimate blur kernels, because blur kernel estimation using accelerometer measurements requires estimating the depth of a scene, the device orientation to compensate the gravity effect, and the initial velocity of the device, which is often error-prone in practice. Moreover, rotational motion can be assumed as the dominant factor of camera shake blur as discussed by Whyte et al. [48].
In the deep learning era, there have been a few works that exploit gyro data. Mustaniemi et al. [30] propose a DNN-based approach that uses gyro data for the first time. They introduce a U-Net-based architecture [36] that takes a concatenation of a blurred image and the gyro data as input, which is also adopted by Lee et al. later [25]. Ji et al. [16] present a deformable-convolution-based network module for handling gyro data, which computes deformable convolution offsets from input gyro data to transform convolution filters according to the blur direction. However, these methods assume that the gyro data is accurate, neglecting potential errors. Moreover, they assume smooth camera motions and model camera motions using a motion vector field consisting of either one [30] or two motion vectors [16] per pixel, or a couple of homographies [25] representing the beginning and end position of the camera during the exposure time. Unfortunately, such restrictive assumptions do not hold for real-world blurred images with complex camera shakes as discussed in Sec. 1.
Utilizing gyro data by converting it into the form of motion vectors for single image deblurring can be considered as a non-blind deconvolution approach. While non-blind deconvolution methods have been studied for decades [33, 49, 22, 18, 26], they are not robust to blur kernel errors and produce unsatisfactory results with erroneous blur kernels. To overcome such limitation, several works adopted hand-crafted priors such as sparse prior [15], hyper-Laplacian prior [46, 22] and deep residual prior [43]. However, they may fail to utilize erroneous blur kernels where such prior assumptions do not hold.
@@ 3 GyroDeblurNet
The goal of GyroDeblurNet is to estimate a sharp deblurred image $S$ from an input blurred image $B$ and its corresponding gyro data $G$ possibly with large errors. Specifically, we define $G$ as a sequence of gyro data such that $G=\left\{g_{0},\cdots,g_{T-1}\right\}$, where $g_{t}$ is the $t$-th gyro data consisting of three angular velocities during the exposure time. $T$ is the number of gyro data, which is proportional to the exposure time. For example, $T=20$ if the exposure time is 0.1 seconds, and the sampling rate of the gyro sensor is 200 Hz. Instead of directly using $G$, we devise a novel gyro data embedding scheme named camera motion field to effectively handle complex camera shakes of an arbitrary exposure duration. In the rest of this section, we explain the gyro data embedding scheme based on the camera motion field, network architecture, and training strategy of GyroDeblurNet in detail.
@@
Figure 3: Detailed architectures of the gyro refinement block and gyro deblurring block.
@@ 3.1 Gyro Data Embedding
GyroDeblurNet first converts input gyro data sequence $G$ into a camera motion field $\mathcal{V}$ before feeding it into the subsequent neural network to handle temporally complex camera shakes of an arbitrary length $T$. The conversion process is illustrated in Fig. 1. Our camera motion field has a fixed channel size $2M$ where $M$ is a hyperparameter, regardless of the length of the input gyro data to allow it to be fed to convolutional neural networks. To achieve this, we first resample the input gyro data sequence $G$ of an arbitrary length $T$ to $M+1$ samples using cubic-spline interpolation. We denote the resampled sequence as $G^{\prime}=\left\{g^{\prime}_{0},\cdots,g^{\prime}_{M}\right\}$ where $g^{\prime}_{m}$ is a resampled gyro sample. The hyperparameter $M$ is set to an even number so that we have the same number of gyro data before and after the temporal center of the exposure duration.
@@
Then, assuming that the camera and the gyro shares the same center and there exists neither off-center rotational nor translational motions, we integrate the angular velocities and obtain $M+1$ camera orientations $\theta_{m}$, where $m=0,\cdots,M$, including the camera orientations at the beginning and end of the exposure. Note that, despite our restrictive assumption on the camera motion, GyroDeblurNet can successfully perform deblurring thanks to its robustness to gyro errors. For each $\theta_{m}$, we compute a homography $H_{m}$ as $H_{m}=KR(\theta_{m})K^{-1}$ where $K$ is the camera intrinsic matrix and $R(\theta_{m})$ is the rotation matrix corresponding to $\theta_{m}$. Based on the camera orientation that we chose as a reference, there can be multiple sets of homography $H_{m}$ each of which leads to shifted vector field. To prevent unnecessary shift after deblurring, we assume $\theta_{M/2}$ at the temporal center to be $(0,0,0)$ so that $H_{M/2}$ is an identity matrix when computing $\theta_{m}$.
@@
Finally, from the obtained homographies, we compute a camera motion field $\mathcal{V}$, which illustrates how each pixel is blurred. Specifically, we define a camera motion field $\mathcal{V}$ as a tensor of size $W/{s}\times H/{s}\times 2M$ where $W$ and $H$ are the width and height of the input blurred image, and $s$ is a scaling factor. Then, at each spatial position $(x,y)$ in $\mathcal{V}$, we compute the warped coordinates of $(x,y)$ by $H_{m}$ for all $m$ and compute the difference between temporally consecutive coordinates to obtain $M$ motion guidance vectors at each position. By stacking the vectors for all the spatial positions, we construct $\mathcal{V}$. In our experiments, we set $M=8$ and $s=2$.
@@
Our camera motion field is motivated by the blur field proposed by Mustaniemi et al. [30] and the motion guidance vector (MGV) field introduced by Ji et al. [16]. Unlike previous methods that rely on only one or two vectors to represent camera shakes at each pixel, our camera motion field offers a more sophisticated representation. By adjusting the hyperparameter $M$, it can effectively capture temporally intricate camera shake patterns. Furthermore, our approach leverages the spatial smoothness of camera shakes, utilizing the hyperparameter $s$. As a result, it can provide a detailed representation of complex motions while maintaining a comparable memory footprint.
@@ 3.2 Network Architecture
The network architecture of GyroDeblurNet is illustrated in Fig. 2. GyroDeblurNet takes a blurred image $B$ and a camera motion field $\mathcal{V}$ as input and estimates a deblurred image $S$. The network consists of two modules: an image deblurring module and a gyro module. The image deblurring module takes a blurred image $B$ and performs deblurring with the aid of gyro data. It then produces a residual $R$ that is added back to $B$ to produce $S$. The image deblurring module adopts a U-Net architecture [36] and the NAFBlock [4] as its basic building block. The image deblurring module adopts gyro deblurring blocks in its bottleneck to perform deblurring using gyro features from the gyro module.
On the other hand, the gyro module takes a camera motion field $\mathcal{V}$ and refines it with the aid of image features from the image deblurring module. The gyro module consists of a convolution layer to embed an input camera motion field $\mathcal{V}$ into the feature space, and gyro refinement blocks. The gyro refinement blocks refine gyro features with the aid of image features from the encoder of the image deblurring module. The gyro module progressively refines gyro features through gyro refinement blocks and strided convolution layers, and obtains a gyro feature of the same spatial size as the feature maps of the bottleneck in the image deblurring module. In the following, the gyro refinement block and the gyro deblurring block, which are the two key components of GyroDeblurNet, are explained in detail. For a detailed network architecture including feature dimensions, refer to the supplementary material.
Figure 4: The input camera motion field passes a convolutional layer to produce differently refined camera motion candidates in different channels. The candidates are then weighted by computing similarity with the blur information of the blurred image.
@@
The structure of the gyro refinement block is illustrated in Fig. 3. The gyro refinement block takes a gyro feature computed from the input camera motion field $\mathcal{V}$ as input. Then, it refines the gyro feature using the channel attention mechanism [13] with the aid of an image feature from the encoder of the image deblurring module. Specifically, the gyro refinement block takes an image feature from the image deblurring module as an additional input, and concatenates it with the input gyro feature. It then computes channel-wise attention weights through global average pooling and $1\times 1$ convolution layers, and refines the gyro feature using the obtained channel-wise attention weights. Finally, it applies a convolution layer to further refine the gyro feature.
The intuition behind the refinement process by the gyro refinement block is illustrated in Fig. 4. The camera motion information in the input camera motion field $\mathcal{V}$ may contain a significant amount of error. From such an erroneous camera motion information, the convolution layer at the beginning of the gyro module produces a gyro feature consisting of multiple camera motion candidates at different channels by introducing different perturbations to the input camera motion information. Then, the gyro refinement block adaptively selects the channels of the gyro feature that are consistent with the blur information encoded in the image feature using the channel attention weights computed from both gyro and image features [13].
@@
The structure of the gyro deblurring block is shown in Fig. 3. The gyro deblurring block performs deblurring of the image feature with the gyro feature from the gyro module. The input image $B$ may have spatially-variant blur caused by camera shakes and moving objects. To effectively remove such spatially-varying blur in the image feature, the gyro deblurring block adopts the deformable convolution operation [7, 55], whose offsets are computed based on both image and gyro features. As the gyro feature provides accurately refined information on the camera shakes, and the image feature provides information on the blur in $B$ including object motion blur, the gyro deblurring block can successfully handle both camera shakes, and object motion blur.
To further remove the remaining spatially-variant blur in the deblurred image feature, the gyro deblurring block applies a spatial attention block to the deblurred image feature following Xu et al. [50]. The deblurred feature is then further refined through a NAFBlock and a convolution layer. While deformable convolution has also been employed by EggNet [16] to deblur images using gyro data, EggNet applies it without accounting for real-world gyro errors. This oversight leads to subpar results for real-world images, as will be shown in Sec. 5. In contrast, our approach incorporates network modules, a training strategy, and data specifically designed to accommodate real-world gyro errors.

### v3

@@ Abstract
In this paper, we present GyroDeblurNet, a novel single-image deblurring method that utilizes a gyro sensor to resolve the ill-posedness of image deblurring. The gyro sensor provides valuable information about camera motion that can improve deblurring quality. However, exploiting real-world gyro data is challenging due to errors from various sources. To handle these errors, GyroDeblurNet is equipped with two novel neural network blocks: a gyro refinement block and a gyro deblurring block. The gyro refinement block refines the erroneous gyro data using the blur information from the input image. The gyro deblurring block removes blur from the input image using the refined gyro data and further compensates for gyro error by leveraging the blur information from the input image. For training a neural network with erroneous gyro data, we propose a training strategy based on the curriculum learning. We also introduce a novel gyro data embedding scheme to represent real-world intricate camera shakes. Finally, we present both synthetic and real-world datasets for training and evaluating gyro-based single image deblurring. Our experiments demonstrate that our approach achieves state-of-the-art deblurring quality by effectively utilizing erroneous gyro data.
@@ 1 Introduction
Blur caused by camera shakes severely degrades image quality as well as the performance of various computer vision tasks. To overcome the image degradation caused by blur, single image deblurring has been extensively studied for decades [5, 11, 53, 33, 49, 16, 14]. Recently, various deep neural network (DNN)-based approaches have been proposed and have shown significant performance improvements over classical approaches. [31, 54, 44, 21, 22, 52, 6, 3]. Nonetheless, they still fail on images with large blur due to the severe ill-posedness of deblurring.
@@
To address the ill-posedness of image deblurring, several attempts have been made to exploit the gyro sensors that most smartphones are now equipped with. Classical approaches based on blur models use gyro data to guide blur kernel estimation [16, 41, 42, 10, 14, 29]. However, their performance is limited by restrictive blur models that struggle to handle noise, nonlinear camera response functions, saturated pixels, moving objects, and other challenges effectively. With the advancement of deep learning, DNN-based approaches have also been proposed to exploit gyro data [30, 23, 15, 34, 46]. Thanks to the remarkable capabilities of DNNs and the rich information from gyro sensors, these methods demonstrate superior performance compared to the classical gyro-based approaches.
Figure 1: Our method shows robust performance even with errors in real-world gyro data. (a) A real-world blurry image. (b) Gyro data visualization onto the blurry image. The gyro data do not perfectly match the blur trajectories due to errors. (c) Our deblurred result.
@@
However, recent DNN-based methods that utilize gyro data are still limited in handling real-world blurred images. First, gyro data from mobile devices such as smartphones usually contain a significant amount of noise caused by various sources [17, 8, 16]. Second, gyro sensors measure the angular velocity around their center. Therefore, recent gyro-based methods assume that the gyro sensor and the camera share the same center and that the camera is only affected by rotational motions. However, in practice, the positions of the camera and the gyro sensor differ, and the camera experiences more complex shakes. Third, real-world blurred images may have moving objects with different blur trajectories that cannot be captured by gyro sensors. All these factors cause the motion information encoded in the gyro data to be inconsistent with the blur in the blurred image (Fig. 1-(b)), making it challenging to exploit gyro data in real-world image deblurring. We refer to such inconsistency between the gyro data and blur as gyro error.
@@
In this paper, we propose GyroDeblurNet, a novel gyro-based single image deblurring approach that produces high-quality deblurring results even in the presence of gyro error (Fig. 1-(c)). To address gyro error, GyroDeblurNet adopts a carefully designed network architecture that includes novel gyro refinement blocks and gyro deblurring blocks. The gyro refinement block corrects gyro data errors using the blur information from the input image. On the other hand, the gyro deblurring block removes blur from the input image using the refined gyro data. Additionally, to further compensate for gyro error, the gyro deblurring block leverages both the blur information in the input image and the input gyro data. Despite these blocks, training a neural network with erroneous gyro data is challenging, as the network might be trained to ignore the input gyro data and rely solely on the input image. To resolve this issue, we propose a curriculum-learning-based training strategy [1].
@@
To handle complex real-world camera shakes, we also present the camera motion field, a novel gyro data embedding scheme that can represent complex camera motions. Real-world blurry images often exhibit complex blur trajectories. Nonetheless, recent gyro-based approaches [30, 23, 15, 34] assume smooth camera motions and represent camera motions using one or two motion vectors per pixel or a couple of homographies for the entire image, which leads to low-quality deblurring results. To resolve this issue, our camera motion field is designed to represent complex camera motions using multiple vectors per pixel while maintaining a comparable memory footprint exploiting the spatially smooth nature of camera shakes.
@@
Previous methods [30, 23, 15, 34] rely on synthetic datasets to train and evaluate the methods. However, their gyro data do not reflect real-world camera shakes that occur when capturing still-shot images either. Gyro data of previous methods were generated by random sampling [55] or obtained from the Visual-Inertial dataset [39]. However, the Visual-Inertial dataset was originally developed for visual odometry and SLAM, so its camera motions were collected from constantly moving cameras, which differ from camera shakes of still-shot images.
@@
For training and evaluating gyro-based deblurring for real-world blurred images, we propose two datasets: GyroBlur-Synth and GyroBlur-Real. GyroBlur-Synth is a synthetic dataset for both training and evaluation. The dataset consists of sharp and blurred image pairs with their corresponding gyro data. The gyro data are acquired from a smartphone device to reflect real-world camera shakes. The blurred images are synthetically generated using the acquired gyro data and the blur synthesis process of RSBlur [36] for realistic blurred images. GyroBlur-Real is a real-world blur dataset for qualitative evaluation with no ground-truth sharp images, and it provides real-world blurred images paired with gyro data collected from a smartphone.
@@
  • We propose a novel gyro-based single image deblurring approach, GyroDeblurNet. Extensive quantitative and qualitative evaluations show that GyroDeblurNet outperforms existing deblurring methods.
  • To handle gyro error, we introduce a carefully designed network equipped with novel gyro refinement and gyro deblurring blocks. We also develop a novel gyro data embedding to represent complex camera shakes and propose a curriculum learning-based training scheme.
@@ 2 Related Work
To resolve the ill-posedness of deblurring, there have been attempts to leverage inertial measurement sensors such as the gyro sensor and accelerometer as they provide valuable information on the camera motion. Joshi et al. [16] and Hu et al. [14] exploit the gyro sensor and accelerometer to estimate spatially-varying blur kernels. On the other hand, Park and Levoy [10] and Mustaniemi et al. [29] use only the gyro sensor to estimate blur kernels, because blur kernel estimation using accelerometer measurements requires estimating the depth of a scene, the device orientation to compensate the gravity effect, and the initial velocity of the device, which is often error-prone in practice. Moreover, rotational motion can be assumed as the dominant factor of camera shake blur as discussed by Whyte et al. [49].
In the deep learning era, there have been a few works that exploit gyro data. Mustaniemi et al. [30] propose a DNN-based approach that uses gyro data for the first time. They introduce a U-Net-based architecture [37] that takes a concatenation of a blurred image and the gyro data as input, which is also adopted by Lee et al. [23]. Ji et al. [15] adopt deformable convolutions to adjust convolution filters based on gyro data. Ren et al. [34] propose a transformer-based network that utilizes gyro data. However, these methods assume accurate gyro data, neglecting potential errors. Moreover, they assume smooth camera motions and model camera motions using a vector field consisting of either one [30, 34] or two vectors [15] per pixel, or a couple of homographies [23] representing the beginning and end position of the camera during the exposure time. Unfortunately, such restrictive assumptions do not hold in real-world cases as discussed in Sec. 1.
Deblurring using gyro data can be framed as non-blind deconvolution, with gyro data serving as blur kernels. However, most non-blind deconvolution approaches assume accurate blur kernels without errors, making them unsuitable for erroneous gyro data [35, 50, 24, 20]. Some attempts have been made to address kernel errors in non-blind deconvolution. Vasu et al. [47] apply varying regularization strengths in a non-blind deblurring algorithm and use a neural network to combine results and refine artifacts induced by kernel errors. Nan et al. [32] tackle kernel error artifacts by estimating residual errors induced by kernel errors using a CNN. However, these approaches assume mild errors in kernels and fail to handle the significant errors presented in gyro data, as will be demonstrated in our experiments.
@@ 3 GyroDeblurNet
The goal of GyroDeblurNet is to estimate a sharp deblurred image $D$ from an input blurred image $B$ and its corresponding gyro data $G$ possibly with large errors. Specifically, we define $G$ as a sequence of gyro data such that $G=\left\{g_{0},\cdots,g_{T-1}\right\}$, where $g_{t}$ is the $t$-th gyro data consisting of three angular velocities during the exposure time. $T$ is the number of gyro data, which is proportional to the exposure time. Instead of directly using $G$, we devise a novel gyro data embedding scheme named camera motion field to effectively handle complex camera shakes of an arbitrary exposure duration. In the rest of this section, we explain the gyro data embedding scheme, network architecture, and training strategy of GyroDeblurNet in detail.
@@
Figure 4: Detailed architectures of the modules in the GyroDeblurNet. (a) Architecture of the gyro refinement block. (b) Architecture of the gyro deblurring block.
@@ 3.1 Gyro Data Embedding
GyroDeblurNet first converts input gyro data sequence $G$ into a camera motion field $\mathcal{V}$ to handle temporally complex camera shakes of an arbitrary length $T$. The conversion process is illustrated in Fig. 2. A camera motion field has a fixed channel size $2M$ where $M$ is a hyperparameter representing the number of vectors in camera motion field, regardless of the length of the input gyro data to allow it to be fed to convolutional neural networks. To achieve this, we first resample the input gyro data sequence $G$ of an arbitrary length $T$ to $M+1$ samples using cubic-spline interpolation. We denote the resampled sequence as $G^{\prime}=\left\{g^{\prime}_{0},\cdots,g^{\prime}_{M}\right\}$ where $g^{\prime}_{m}$ is a resampled gyro sample. The hyperparameter $M$ is set to an even number so that we have the same number of gyro data before and after the temporal center.
@@
Then, assuming that the camera and the gyro shares the same center and there exists neither off-center rotational nor translational motions, we integrate the angular velocities and obtain $M+1$ camera orientations $\theta_{m}$, where $m=0,\cdots,M$, including the camera orientations at the beginning and end of the exposure. Note that, despite our restrictive assumption on the camera motion, GyroDeblurNet can successfully perform deblurring thanks to its robustness to gyro errors. For each $\theta_{m}$, we compute a homography $H_{m}$ as $H_{m}=KR(\theta_{m})K^{-1}$ where $K$ is the camera intrinsic matrix and $R(\theta_{m})$ is the rotation matrix corresponding to $\theta_{m}$. To prevent unnecessary shift after deblurring, we chose the temporal center as a reference by assuming $\theta_{M/2}$ at the temporal center to be $(0,0,0)$ so that $H_{M/2}$ is an identity matrix when computing $\theta_{m}$.
@@
Finally, from the obtained homographies, we compute a camera motion field $\mathcal{V}$, which illustrates how each pixel is blurred. Specifically, we define a camera motion field $\mathcal{V}$ as a tensor of size $W/{s}\times H/{s}\times 2M$ where $W$ and $H$ are the width and height of the input blurred image, and $s$ is a scaling factor. Then, at each position $(x,y)$ in $\mathcal{V}$, we compute the warped coordinates of $(x,y)$ by $H_{m}$ for all $m$ and compute the difference between temporally consecutive coordinates to obtain $M$ vectors. By stacking the vectors for all the spatial positions, we construct $\mathcal{V}$. In our experiments, we set $M=8$ and $s=2$. Fig. 1-(b) visualizes camera motion field vectors onto a real-world blurry image.
@@
Note that similar representations for gyro data based on vector fields have also been proposed by previous approaches [30, 15], as discussed in Sec. 2. However, unlike these methods that rely on only one or two vectors to represent camera shakes at each pixel, our camera motion field offers a more sophisticated representation. By adjusting the hyperparameter $M$, it can effectively capture temporally intricate camera shake patterns. Furthermore, our approach leverages the spatial smoothness of camera shakes, utilizing the hyperparameter $s$. As a result, it can provide a detailed representation of complex motions while maintaining a comparable memory footprint.
@@ 3.2 Network Architecture
GyroDeblurNet takes a blurred image $B$ and a camera motion field $\mathcal{V}$ as input and estimates a deblurred image $D$. The network consists of two modules: an image deblurring module and a gyro module (Fig. 3). The image deblurring module deblurs a blurred image $B$ with the aid of gyro data. It then produces a residual $R$ that is added back to $B$ to produce $D$. The image deblurring module adopts a U-Net architecture [37] and the NAFBlock [3] as its basic building block. The image deblurring module adopts gyro deblurring blocks in its bottleneck to perform deblurring using gyro features from the gyro network.
On the other hand, the gyro module takes a camera motion field $\mathcal{V}$ and refines it with the aid of image features from the image deblurring module. The gyro module consists of a convolution layer to embed an input camera motion field $\mathcal{V}$ into the feature space, gyro refinement blocks and strided convolution layers. The gyro refinement blocks refine gyro features with the aid of image features from the encoder of the image deblurring module. The gyro module progressively refines gyro features through gyro refinement blocks and strided convolution layers, and obtains a gyro feature of the same spatial size as the feature maps of the bottleneck in the image deblurring module. In the following, the gyro refinement block and the gyro deblurring block, which are the two key components of GyroDeblurNet, are explained in detail. For a detailed network architecture including feature dimensions, refer to the supplementary material.
@@
Fig. 4-(a) illustrates the structure of the gyro refinement block. The camera motion information in the input camera motion field $\mathcal{V}$ may contain significant global errors caused by gyro sensor noise, unknown rotational centers, and missing translational motions, as visualized in Fig. 1-(b). The gyro refinement block is designed to compensate for these global errors using the blur information in the input image based on the following intuition.
The gyro refinement block takes a gyro feature computed from the input camera motion field $\mathcal{V}$ as input. This input gyro feature encodes multiple camera motion candidates across different channels, obtained by introducing perturbations to the input camera motion information in the previous layers. To globally refine the gyro feature, the gyro refinement block adaptively selects the channels of the gyro feature that are consistent with the blur information in the input image. Specifically, the gyro refinement block takes an image feature from the image deblurring module as an additional input and concatenates it with the input gyro feature. It then computes channel-wise weights [13] through global average pooling and $1\times 1$ convolution layers, and multiplies them to the gyro feature. Following the gyro refinement block, a strided convolution layer further refines the gyro feature while reducing its spatial resolution.
@@
The gyro deblurring block, illustrated in Fig. 4-(b), performs deblurring of the image feature using the gyro feature from the gyro module. The input image $B$ may have spatially-variant blur caused by camera shakes and moving objects. To effectively remove such blur, the gyro deblurring block consists of two sub-blocks: the first sub-block performs spatial refinement of the blur information and deblurring of the image feature using the refined blur information, while the second sub-block further refines the deblurred image feature.
Specifically, the first sub-block concatenates the image feature and gyro feature, computing spatially-refined blur information in the form of deformable convolution kernel offsets. It then performs deblurring of the image feature using these offsets. The second sub-block further refines the deblurred image feature using spatial attentions [51]. Spatial attention weights are computed from the deblurred image block through convolution and sigmoid layers, and multiplied with the deblurred image feature, which is further processed through convolution and concatenation layers.